bobo424:

What is the solution to the linear equation? x – StartFraction 2 Over 3 EndFraction x minus StartFraction one-half EndFraction equals StartFraction 1 Over 3 EndFraction plus StartFraction 5 Over 6 EndFraction x. = + x

Mercury:

x - (2/3)x - (1/2) = (1/3) + (5/6)x ? if I'm transcribing this correctly? if so, you can combine like terms. subtracting (5/6)x from both sides will give you x - (2/3)x - (5/6)x - (1/2) = (1/3) adding 1/2 to both sides x - (2/3)x - (5/6)x = (1/3) + (1/2) from there it's a matter of finding common denominators so you can combine the fractions. on the left, you have denominators 1, 3, and 6 (remember that 1 = 1/1) so find the least common multiple of these numbers and make that the new denominator through multiplication. similar logic with the right side.